The volume of a box V, varies with some variable x as V(x)=x^3 - 12x ^2 + 44x -48 cubic metres. If (x - a) metre is the measurement of one side of the box, then find the value for a.



1
Expert's answer
2021-10-04T04:58:24-0400
"V(x)=x^3 - 12x ^2 + 44x -48"

"x^3 - 12x ^2 + 44x -48 =x^2(x-2)-10x(x-2)+24(x-2)"

"=x^2(x-2)-10x(x-2)+24(x-2)"

"=(x-2)(x^2-10x+24)"

"=(x-2)(x-4)(x-6)"

"a_1=2, a_2=4, a_3=6"


"a\in\{2,4,6\}"